OpenStudy (anonymous):

How can you tell when a quadratic equation has two identical, rational solutions? when the radicand is negative when b in the quadratic formula is greater than the radicand when the radicand equals zero when the radicand is not a perfect square

OpenStudy (anonymous):

when radicand = zero

OpenStudy (anonymous):

i assume that radicand = b^2-4ac ??

OpenStudy (anonymous):

in uk we call it the discriminant

OpenStudy (anonymous):

when the radicand equals zero for ax^2+bx+c=0 find b^2-4ac if that is = 0 ------> the equation has 2 identical roots

OpenStudy (zarkon):

it is called discriminant in the US also. In general, the radicand is the expression under a radical sign.